OpenStudy (anonymous):

Find the differential equation of families of curve: All circles. Use the curvature

OpenStudy (anonymous):

first write the general eq. of circle . Let (h,k) be the centre and r be the radius of the circle eq. of circle is\[\left( x-h \right)^{2}+\left( y-k \right)^{2}=r ^{2}\] diff. twice w.r.t x to eliminate a & b \[2\left( x-h \right)+2\left( y-k \right)y'=0\] \[or \left( x-h\right)+\left( y-k \right)y'=0\] \[or \left( x-h \right)=-\left( y-k \right)y'....(1)\] again diff. 1=-[y'.y'+((y-k)y"] \[or 1=-[\left( y' \right)^{2}+\left( y-k \right)y"]\] \[\left( y-k \right)y"=-\left( 1+\left( y' \right)^{2}\right)\] \[(y-k)=-\frac{ 1+\left( y' \right)^{2} }{ y" }\]

OpenStudy (anonymous):

from (1) \[ x-h=+\frac{ 1+\left( y' \right) ^{2}}{y" }y'...(2)\]

OpenStudy (anonymous):

plug the values of x-h and y-k in \[\left( x-h \right)^{2}+\left( y-k \right)^{2}=r ^{2}\] \[solve and put y'=\frac{ dy }{dx },y"=\frac{ d ^{2}y }{ dx ^{2} }\]

OpenStudy (anonymous):

i am not sure you want like this or something else.

OpenStudy (anonymous):

can u give the final answer?

OpenStudy (anonymous):

\[\left[ \frac{ 1+\left( y' \right)^{2} }{y"} \right]^{2}\left( y' \right)^{2}+\left[ \frac{ 1+\left( y' \right)^{2} }{ y" } \right]^{2} =r ^{2} \] \[\left\{ 1+\left( y' \right)^{2} \right\}^{2}\left\{ \left( y' \right)^{2} +1\right\}=r ^{2 }\left( y" \right)^{2}\] \[\left[ 1+\left( y' \right)^{2} \right]^{3}=r ^{2}\left( y" \right)^{2}\] \[\left[ 1+\left( \frac{ dy }{ dx } \right)^{2} \right]^{3}=r ^{2}\left( \frac{ d ^{2} y}{dx ^{2} } \right)^{2}\]